(10, -1) and (0, 0). What is its equation in slope-intercept form?
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

(10, -1) and (0, 0). What is its equation in slope-intercept form?
----------------
Find the slope, m
m = diffy/diffx
====================
Then use y-y1 = m*(x-x1) where (x1,y1) is either point.
============================================
This is basic and elementary.
You have know this as you know 2+2 = 4


Answer by ikleyn(52944)

.

On this subject,  look into the lesson
    - Equation for a straight line in a coordinate plane passing through two given points
in this site.

Read it and then solve your problem on your own.

From this lesson,  learn the subject once for all.